Status Report

NOAA GOES SATOPS Morning Report 25 Jan 2001

By SpaceRef Editor
January 25, 2001
Filed under ,

SATOPS MORNING REPORT January 25, 2001

POES

NOAA-16 rev 01764 / F at 1220z on January 24: Unable to receive High
Rate Picture Transmission (HPRT) during the pass.  Beacon Telemetry
was used to provide TIROS Information Processor (TIP) data and pass was
otherwise nominal.  Post-pass investigation found the problem to be
SPFS2 having “transitioned into an unknown state”, which caused it to fail
to lock on incoming HRPT.  It was reset by powering it off and on,
after which it performed nominally.

At 1315z on January 24: Frame Sync anomaly interrupted wideband comm
from Fairbanks and resulted in a halt to the Frame Sync test that started
on January 23.  The test plan was amended immediately and testing
now continues.

 

DMSP

No change in the status of the DMSP spacecraft.

DMSP operations were nominal over the past 24 hours.

 

GOES

No change in the status of the GOES spacecraft.

At 1527z on January 24: GOES-8 image frame break due to loss of signal
synch.

 

 

Questions or comments to [email protected]

SpaceRef staff editor.